The following transactions occurred during January, the first month of operations for Aly Services, Inc. (ASI):

a. Toby and Aly incorporate ASI, and each owner contributes $10,000 cash $20,000 total) to the new corporation. The corporation issues stock to Toby and Aly b. ASI Borrows $5,000 cash from a local bank using a one-year note c. ASI pays $10,000 cash to buy a truck d. ASI purchases $400 worth of grooming supplies (shampoos, etc.) on account. The company receives a bill and will pay the bill later e. On January 1, ASI pays $600 in advance to Geico to cover insurance for ts truck from January 1 through June 30 f. ASI groomed 30 dogs at a rate of $40 per dog ($1,200 total) during the month. The company collected cash from clients immediately after providing the service g. ASI has a contract with two large corporate clients where ASI cares for all of the clients employees pets for the month while the employees are at work On January 15, ASI provided $6,000 worth of dog sitting and grooming services and gave each of the two clients a bill for the services. The clients did not pay ASI any cash yet. h. On January 28, a corporate client purchased a 12-month unlimited package and immediately paid ASI $24,000 cash. ASI will provide the client with grooming services over the next 12 months, beginning orn February 1, and ending on January 30 of next year i. During January ASI paid $1,500 Cash to Employees for work performed j. One of the corporate clients from transaction (g) paid their bil by giving k. ASI pays $250 to their supplier for the grooming supplies purchased at I. ASI pays a $2,000 dividend to its shareholders in January ASI $4,000 cash transaction (d)Prepare an unadjusted trial balance for ASI, Inc. for the month ended January 31 ASI, INC Unadjusted Trial Balance Account Name Debit Credit Totals
Answer #1

BALANCES :-

Debtor= 6000-4000=2000

Creditor=400-250=150